OpenStudy (anonymous):

evaluate xy^3-xy if x=3 and y= -2 a. 18 b. 30 c. -18 d. 4 e. none of the above

OpenStudy (anonymous):

I'm getting B. 30 ?

OpenStudy (anonymous):

what symbol is ^

OpenStudy (aum):

Can you show the work how you got 30?

OpenStudy (anonymous):

Yes, this is how i did it:

OpenStudy (anonymous):

is this a t?

OpenStudy (aum):

xy^3-xy if x=3 and y= -2 (3) * (-2)^3 - (3) * (-2) = 3 * (-8) - (-6) = -24 + 6 = -18

OpenStudy (anonymous):

whoops i didn't realize it was a ^3, i was doing it with a ^2 i'm gonna redo it now and see what i get

OpenStudy (anonymous):

i also got 210

OpenStudy (anonymous):

please tell me what ^ means

OpenStudy (anonymous):